OpenStudy (anonymous):

I need help with trigonometric substitions Intergral of (sqrt(y^2-49)/y)dy I know that y will be equal to 7sec theta My solutions manual has a funny answer that I can't get to. I need help working it out, thanks!

OpenStudy (anonymous):

\[\int\limits (\sqrt{y^2-49})/y) dy\]

myininaya (myininaya):

\[\int\limits_{}^{}\frac{\sqrt{49\tan^2{\theta}-49}}{7\tan \theta} 7 \sec^2{\theta} d \theta\] \[\frac{\sqrt{49}*7}{7}\int\limits_{}^{}\frac{\sqrt{\tan^2{\theta}-1}}{\tan \theta} \sec^2{\theta} d \theta\]

myininaya (myininaya):

so we can conclude that tan was the wrong substituition

OpenStudy (anonymous):

My solutions manual has this: y=7sec\[\theta\] dy=7sec\[\theta\]tan\[\theta\]d\[\theta\]

OpenStudy (anonymous):

oh darn...i wonder why it spaces it like that

OpenStudy (anonymous):

y=7sec theta dy=7sec theta tan theta and therefore \[\sqrt{(y^2-49)}\] is equal to 7tan theta

OpenStudy (anonymous):

but i have no clue how they got 7 tan theta for the \[\sqrt{y^2-49}\]

myininaya (myininaya):

oops thats wrong i cant think today i sneezed so much

OpenStudy (anonymous):

Haha, I was all confused, lol!

OpenStudy (anonymous):

It's all good tho

myininaya (myininaya):

ok im going to write this on paper and post it k? give me a few

OpenStudy (anonymous):

Oh thanks!

myininaya (myininaya):

\[7\int\limits_{}^{}\tan^2{\theta} d \theta=7\int\limits_{}^{}(\sec^2{\theta}-1) d \theta=7\int\limits_{}^{}\sec^2{\theta} d \theta-7\int\limits_{}^{}1 d \theta\] =\[7\tan \theta-7 \theta+C\] we need to put this in terms of x

myininaya (myininaya):

or i mean in terms of y lol do you see our traingle? \[\tan \theta=\frac{opposite}{adjacent}=\frac{\sqrt{y^2-49}}{7}\]

myininaya (myininaya):

\[\sec{\theta}=\frac{y}{7}\] so \[\theta=\sec^{-1}(\frac{y}{7})\]

OpenStudy (anonymous):

Yes! Thank you so much! Duh, I can't believe I didn't see that!

myininaya (myininaya):

\[7*\frac{\sqrt{y^2-49}}{7}-7*\sec^{-1}(\frac{y}{7})+C\]

myininaya (myininaya):

those 7 cancel in the first fraction

myininaya (myininaya):

but u knew that

myininaya (myininaya):

it always helps me to draw a traingle for these type of problems

myininaya (myininaya):

thats what i should had done at the very beginining

OpenStudy (anonymous):

Yea, I need to draw the triangle first thing, I always do it last, hehe

myininaya (myininaya):

drawing those three triangles helps me to determine which substituition to use also it helps when you need to put everything back in terms of whatever it was

myininaya (myininaya):

try another one and if you can't get it post it

OpenStudy (anonymous):

Awesome, thanks!